You go to class every MWF. The probability that you will late to a class is 1/5. If you are late on a day, then probability that you will be late again next class is 1/10. What is the probability that you will be late exactly twice in a week when you go to the MWF class?
Answer #1

let L and T are event of being late and on time on any given day

probability that you will be late exactly twice in a week when you go to the MWF class

=P(late on Monday and Wednesday)+P(late on Monday and Friday)+P(late on wednesday and Friday)

=P(LLT)+P(LTL)+P(TLL)

=(1/5)*(1/10)*(9/10)+(1/5)*(9/10)*(1/5)+(4/5)*(1/5)*(1/10)

=7/100 =0.07